You're going to build a 3 foot by 7 foot garden. If you surround it with a 1 foot wide brick border how many square feet of brick do you need?
Answer by ikleyn(52776)

.

The outer dimensions of your construction are  1 + 3 + 1 = 5 ft  and  1 + 7 + 1 = 9 ft,


so the full area is  5*9 = 45 square feet.


To answer the question, subtract from it the area of the garden itself, and you will get 

    
      45 - 3*7 = 45 - 21 = 24 square feet.    ANSWER